Write a balanced equation for the complete combustion of pentane, C5H12. You may OMIT state symbols (s, l, g, aq, etc.). Enter the alkane formula as C5H12.
Answer #1

CxHy+(x+y/4)O2-----> xCO2+(y/2)H2O

here x=5 y=12

x+y/4 = 5+12/4 = 5+3 =8

C5H12+8O2-----> 5CO2+ 6 H2O